# Channel Choice and Customer Value

### v1

@@ Abstract
The rapid growth of digital shopping channels has prompted many traditional retailers to invest in e-commerce websites and mobile apps. While prior literature shows that multichannel customers tend to be more valuable, it overlooks how the reason for adopting a new channel may shape post-adoption behavior. Using transaction-level data from a major Brazilian pet supplies retailer, we examine how adoption of online shopping—by previously offline-only customers—affects post-adoption spend, profitability, and channel usage. We study four distinct adoption pathways: organic adoption, adoption due to the COVID-19 pandemic, Black Friday promotions, and a newly launched loyalty program. We find that although all adopter groups increase their spending relative to offline-only customers, post-adoption behavior differs based on the reason for adoption. COVID-19 adopters behave similarly to organic adopters in terms of spending, but show greater offline stickiness consistent with consumer inertia and habit theory, yielding higher profits due to higher offline margins. In contrast, promotion-driven adopters spend less post-adoption due to forward buying and exhibit lower profitability. Our findings caution against treating all multichannel customers as equal and highlight the importance of incorporating behavioral theories into forecasting and targeting strategies. Firms should account for adoption motives when evaluating channel and promotional investments.
@@ 1 Introduction
Online shopping has been growing in popularity over the last two decades, with the proliferation of computers and smartphones. In response, many grocery and retail stores like Walmart, Target, Costco, etc., which earlier only operated physical retail stores, now sell their products through both physical stores and e-commerce channels such as websites and mobile apps. Given the size of the retail industry and the growth of digital channels (especially the acceleration of the adoption of online shopping during the COVID-19 pandemic), quantifying the effect of the adoption of digital shopping channels is important for managers and firms (Goyal 2023; Bloomberg 2020).
In this paper, we consider two related research questions. First, we examine whether and how the adoption of digital shopping channels (e-commerce website and/or mobile app) of a retailer affects the purchase behavior of consumers who were previously purely offline-only (brick-and-mortar) shoppers. We seek to document the impact of the adoption of online shopping on three metrics of interest for managers: (1) Spend, (2) Profitability, and (3) Stickiness to the offline channels post-adoption. The first metric, spend, quantifies the total revenue/sales (across offline and online channels) generated by the customer after the adoption of online shopping, the second metric quantifies the profits (sales minus cost) generated by the customer, and the third metric denotes the extent to which consumers use offline channels even after adopting online shopping. Second, we are interested in understanding whether and how post-adoption shopping behavior differs depending on consumers’ different reasons for adopting online shopping. This is important from a managerial perspective – if the changes to consumers’ purchase behavior (post-online adoption) are largely driven by the motivations/reasons for consumers choosing to adopt online shopping, then it will be challenging to extrapolate the behavior of organic adopters (who adopt online shopping without any external incentives) to other consumers who are incentivized to adopt by marketing campaigns or other external economic or environmental factors i.e., different motivations compared to organic adopters. Furthermore, if the differences in post-adoption behavior relative to organic adopters depend on the type of external factor (i.e., macroeconomic or environmental vs. firm-driven promotions); forecasting post-adoption behavior would be even more complicated. On the other hand, if the consumers who adopt online shopping due to different reasons behave similarly post-adoption, then managers can be more confident in extrapolating the behavior of those who adopt organically to a broader set of adopters (and be less concerned about the reason for the adoption of online shopping).
While a large stream of prior research suggests that multichannel consumers (defined as consumers who buy from multiple channels including catalogs, mobile apps, stores, websites, etc.) tend to spend more than single-channel consumers (Neslin and Shankar 2009), it is not obvious if the earlier results translate directly to our setting for three reasons. We briefly discuss the reasons here and refer the reader to Table 1 for a more detailed comparison to previous literature. First, they do not specifically study the effect of the adoption of online shopping from previously only purchasing from brick-and-mortar stores. Studying the physical store to online (e-commerce or mobile app) transition is both novel and interesting because many major retailers started out with brick-and-mortar stores, and only over the last couple of decades have been investing more heavily in developing the e-commerce and mobile app channels. E-commerce websites and mobile apps, as channels, are potentially quite different from other channels like mail order catalogs, phone shopping, or interacting with salespeople because both these channels reduce search costs, and it is easy for customers using these channels to visit other retailers’ websites and apps to comparison shop across retailers. Second, the vast majority of these papers adopt one of two approaches – (a) a cross-sectional analysis (Kumar and Venkatesan 2005; Thomas and Sullivan 2005; Kushwaha and Shankar 2013) comparing the spend of single-channel shoppers (any single channel, not necessarily just physical stores) vs. multichannel shoppers (which can include physical stores and online channels, but also mail-order catalogs, ordering through phone, salespeople, etc.), or, (b) a within-customer analysis (Venkatesan and Ravishanker 2007; Ansari, Mela, and Neslin 2008) comparing the pre- and post-adoption periods of customers who start using multiple channels from being single channel only (where again, these channels include websites, catalogs, physical stores etc.). However, cross-sectional comparisons cannot answer the question of how customer behavior changes after the adoption of multiple channels, and within-customer analyses require the strong assumption that the entire difference between the before and after periods can be fully attributed to the adoption of multiple channels, without any other time-varying factors. More recently, some studies have used Difference-In-Differences (DID) analyses to study the effects of the adoption of mobile apps (Narang and Shankar 2019; Wang, Malthouse, and Krishnamurthi 2015). However, before the adoption of the app, customers in these studies could have already been using the retailer’s website to shop online. As such, their findings may not generalize to our context, where the transition in focus is from physical stores to online channels. Third, the previous studies only examine customers who have voluntarily/organically chosen to use multiple channels for shopping or have chosen to adopt the firm’s mobile app. They completely ignore different channel adoption pathways and implicitly assume that the benefits of multichannel customers are uniform regardless of why they chose to adopt multiple channels or a particular channel. There are a couple of important theoretical reasons to expect differences between (a) organic adopters and externally incentivised adopters and (b) different types of external incentives i.e. macro-environmental shocks vs. firm driven promotions. Consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009) would predict that adoption driven by external shocks (e.g., macro shocks or firm-driven promotions) may lead to temporary disruptions in consumer behavior (i.e. adopting a new channels and using multiple channels to shop), but without reinforcement, consumers may revert to prior routines (i.e. not use the newly adopted channels much). Next, for promotion-induced adoption in particular, the theory of forward buying and demand distortion may apply. Promotions incentivizing online adoption may also induce forward buying, leading to reduced post-promotion consumption (Neslin, Henderson, and Quelch 1985; Ailawadi and Neslin 1998), in turn leading to lower revenues for the firm post-adoption. Thus, we extend the previous literature by (a) studying a particularly important type of transition from single-channel to multi-channel (i.e. offline-only to offline-plus-online) and (b) by examining the behavior or consumers who adopted online shopping due to different reasons, and challenging the assumption that all multichannel customers are uniformly beneficial and behave the same after adopting a new channel regardless of their reason for adoption.
@@
We use data from a large pet supplies retailer in Brazil with a large network of brick-and-mortar stores and an online presence (website and mobile app) to answer our research questions. Our data consists of all consumer purchases across all channels for the period of 2019-2024. This time period allows us to construct four groups of adopters of online shopping who previously only shopped at the firm’s physical stores – (1) covid adopters who started shopping online during the onset of COVID-19, (2) Black Friday adopters who adopted online shopping due to deep discounts available online during Black Friday, (3) Loyalty Program adopters who adopted online shopping during the launch of a loyalty program which had to be activated on the retailer’s mobile app, and (4) organic adopters who started shopping online fully of their own volition, without any external stimuli. In addition, we also consider a base control group of non-adopters or offline-only customers (consumers who did not adopt online shopping during our observation period). We then construct consumer-level monthly panel data to quantify the revenue, profitability, and share of spend in the offline and online channels. Since we have four groups of adopters who chose to adopt for different reasons, in addition to non-adopters, we can address both our research questions – quantify the impact of the adoption of online shopping on consumers’ spend, profitability, and stickiness to offline channels, as well as comment on whether and how the different reasons for adoption affects the post-adoption behavior of adopters.
@@
We now discuss our main findings. First, we find that relative to offline-only customers, adopters of online shopping, have higher spend and profitablility in the post-adoption period. However, the post-adoption behaviors of the different types of adopters diverge significantly from each other across the three key dimensions: total spend, profitability and the share of offline/online spending. While covid adopters spend similarly relative to organic adopters post-adoption, both promotion-driven adopters, Black Friday adopters and Loyalty Program adopters, spend less. This reduced spend is driven largely by forward buying during the promotional period, meaning they stocked up and then reduced future purchases, which doesn’t occur in the case of covid adopters. In terms of the share of offline/online spend post-adoption, consistent with consumer inertia and habit theory, all three adopter groups motivated by external stimuli (covid adopters, Black Friday adopters and Loyalty Program adopters) have a lower online spend share post-adoption relative to organic adopters. A combination of demographic characteristics and consumer inertia and habit theory also explain differences in how this share evolves over time. We find that covid adopters show a slower growth in online spend share over time. This is mainly driven by the fact that covid adopters are older and thus are more set in their habits. Black Friday adopters on the other hand, are younger and show a faster growth rate in the online spend share relative to organic adopters, although they still take time to catch up to the same levels. There is no difference in trends for the online spend share between Loyalty Program adopters and organic adopters. This may be due to the fact that the two adopter groups are quite similar in terms of demographics, and also the fact that the Loyalty Program involves monthly online activation, and therefore there is some reinforcement for the habit of online shopping for Loyalty Program adopters. The post-adoption profits for the different adopter groups are a function of both spend and the share of offline/online share. We find that despite similar spend, covid adopters are more profitable post-adoption, than organic adopters in our setting. This is because of their higher spend share in the offline channel, where profit margins are higher in our setting. In contrast, Black Friday adopters and Loyalty Program adopters are less profitable due to their lower overall spending in the post-adoption period.
Our findings suggest that the reason for a consumer’s decision to adopt online shopping significantly shapes their long-term behavior. Consumers who adopt online shopping organically behave very differently from those motivated by external events or firm-led promotions, due to a combination of their demographics, consumer inertia, and forward buying during promotions. Our findings have important managerial implications. For example, had the retailer assumed that Black Friday adopters would behave like organic adopters, it would have significantly underestimated the payback period for discounts offered during the promotion. While projections based on organic adopters would suggest a break-even point at 13 months, the actual break-even occurred only after 39 months. Even when overall spending appears similar—as in the case of covid adopters—differences in channel usage (online vs. offline) can have substantial implications for profitability. Firms should avoid using organic adopters as the benchmark when predicting the behavior of customers adopting a new channel for other reasons. More broadly, our empirical results indicate that managers can use their knowledge of expected customer demographics, and apply the theories of consumer inertia and forward buying to forecast spend, profit and channel usage outcomes in other channel adoption scenarios. Tailoring strategies to the unique characteristics of each type of adopter group is essential for maximizing both customer value and long-term profitability.
In summary, our paper makes both substantive and managerial contributions to the literature on channel choice and consumer value. From a substantive perspective, we show that the adoption of online shopping from only using physical stores, either voluntarily or prompted by external factors like a macro-environmental shock or firm-driven promotions, causes consumers to increase their total spend with the firm relative to consumers who stay offline-only. More interestingly, we show that consumers who adopt online shopping due to different reasons, have different post-adoption outcomes in terms of their spend, channel usage and profits, and these differences can be explained using demographics, consumer inertia theory, and forward buying during promotions. From a managerial perspective, we not only quantify the impact of the adoption of online shopping on revenues and profits but also provide a cautionary tale against naively extrapolating the profitability of one group of online shopping adopters to a broader/different set of adopters. We also provide theoretical explanations of why post-adoption behaviors of different adopter groups differ, which managers can use to forecast outcomes in other settings.
@@ 2 Related Literature
Paper Setting Channels Methodology Type of adopters
@@
(subscribers only) SMS, Internet, external stimulus
@@
There is a large empirical literature in marketing studying the effects of multichannel shopping vs. single-channel shopping. We provide a brief overview of the most closely related papers here in Table 1 and discuss how our paper differs from and contributes to this literature in the rest of this section. We refer readers to Neslin et al. (2006); Neslin and Shankar (2009); Verhoef (2012) for more comprehensive reviews of the broader multichannel literature.
As shown in Table 1, our paper differs from the earlier literature both substantively and methodologically. First, from a substantive perspective, we study a specific type of single-channel to multichannel transition that has not been studied before – from shopping at physical stores only to shopping both at physical stores and online through the retailer’s website/mobile app. Earlier findings from the multichannel literature largely focus on the transition from any single channel to some combination of multiple channels (including physical stores, catalogs, phones, mail orders, salespeople, mobile apps, and websites).
@@
Studying the physical store to e-commerce transition is of particular interest to traditional brick-and-mortar retailers (Goyal 2023; Bloomberg 2020) like Walmart, Target, and Costco since these channels (websites and mobile apps) represent a key area of growth for them. However, since the earlier literature does not explicitly study this transition (and rather considers a broader set of channels), those findings may not generalize to this context. Indeed, there are some important differences among these different channels. For instance, websites and apps provide easy search and comparison between products within a retailer and also between competing retailers because it is easy to jump between retailer websites/apps on your device. This is not the case for other channels that have been studied before, like catalogs, salespersons, or phone shopping. Also, in many cases, prices on a retailer’s website or mobile app are lower than the prices in physical stores, which is not the case for other channels. Thus, the findings from other channels may not fully translate to the channels we study here.

### v2

# Channel Adoption Pathways and Post-Adoption Behavior

@@ Abstract
The rapid growth of digital shopping channels has prompted many traditional retailers to invest in e-commerce websites and mobile apps. While prior literature shows that multichannel customers are more valuable, it overlooks how the motive for adopting a new channel shapes post-adoption behavior. Using transaction-level data from a major Brazilian pet supplies retailer, we study offline-only consumers who adopt online shopping via four distinct pathways: organic adoption, the COVID-19 pandemic, Black Friday promotions, and a loyalty program. We examine how these pathways affect post-adoption spend, profitability, and channel usage using consumer-level panel data and difference-in-differences estimates. We find that all adopters increase spending relative to offline-only consumers, but their post-adoption behaviors differ systematically by adoption motive. Promotion-driven adopters engage in forward buying and exhibit lower subsequent profitability, whereas COVID-19 adopters display stronger offline persistence consistent with consumer inertia and habit theory. Our findings have important managerial implications: firms should design promotions that discourage stockpiling, reinforce habits among customers pushed online by external shocks, and explicitly account for heterogeneity in channel adoption motives when forecasting customer lifetime value and assessing the breakeven and ROI of promotions designed to induce the adoption of new channels.
@@ 1 Introduction
Online shopping has grown rapidly over the last two decades with the proliferation of computers and smartphones. In response, many grocery and retail chains (e.g., Walmart, Target, Costco), historically brick-and-mortar, now also sell through websites and mobile apps. Given the size of retail and the rise of digital channels, it is important to quantify how adopting digital shopping channels affects consumers’ behavior (Goyal 2023; Bloomberg 2020).
Prior research has found that multichannel consumers (those who buy using multiple channels of a retailer) tend to spend more than single-channel consumers; see $\S$2. However, this work largely ignores why consumers adopt new channels, implicitly assuming that the benefits from multichannel adoption are the same regardless of the reasons for adoption of new channels. Yet consumers may adopt a retailer’s online shopping channels (from previously being offline-only) for many different reasons, and it is unlikely that channel adoption pathways have no impact on post-adoption behavior. We classify potential online adoption pathways into three categories: (1) macro-environmental shocks such as COVID-19, nearby store closures, or severe weather events that push consumers online; (2) firm-driven promotions that require engagement with online channels to obtain benefits – this can include both short-term promotional events and holiday sales as well as long-term offers such as loyalty programs; and (3) an organic pathway, in which consumers adopt online shopping of their own volition due to the convenience of an additional channel.
It is important for managers to understand heterogeneity in post-adoption behavior across these pathways. If changes in purchase behavior after online adoption depend on the motivations for adoption, then it is difficult to extrapolate the behavior of consumers who adopt organically (without external incentives) to those induced to adopt through marketing campaigns or external factors. Moreover, if differences relative to organic adopters vary by the type of external factor (e.g., macro shock vs. promotion), forecasting post-adoption behavior becomes even more complicated. Conversely, if consumers who adopt online shopping for different reasons behave similarly post-adoption, managers can extrapolate from organic adopters and worry less about the underlying reason for adoption. These considerations motivate our research questions.
Research Questions and Theoretical Predictions
We consider two research questions. First, how does the adoption of a retailer’s digital shopping channels (e-commerce website and/or mobile app) affect the purchase behavior of customers who were previously purely offline-only (brick-and-mortar) shoppers? We focus on three managerially salient outcomes: (1) total spend, (2) profitability, and (3) stickiness to offline channels post-adoption. Second, conditional on adopting online shopping, how does post-adoption behavior differ across the three adoption pathways (organic, macro-shock-driven, and promotion-driven)?
Building on the multichannel literature and behavioral frameworks (discussed in detail in $\S$2 and $\S$3), we hypothesize that customers who adopt a firm’s online channels and become multichannel shoppers will, on average, increase their total spending relative to offline-only customers. Beyond this average effect, we anticipate substantial heterogeneity across adoption pathways. Consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009) suggest that when adoption is induced by external shocks or incentives rather than intrinsic preference, consumers are more likely to revert toward prior routines once the stimulus subsides. Thus, relative to externally driven adopters, organically adopting customers should exhibit a higher post-adoption share of spending on online channels and faster growth in that share. We further anticipate that such differences will be more pronounced for older, habit-driven customers.
For adoption spurred by time-limited promotions, theories of forward buying and demand distortion (Neslin, Henderson, and Quelch 1985; Hendel and Nevo 2006; Van Heerde and Neslin 2017) imply that promotion-driven adopters may temporarily inflate their spending during the promotional period by stocking up, but then reduce spending afterwards. This mechanism can dampen or even offset the long-run spending uplift typically associated with multichannel adoption. Finally, because profit margins generally differ across online and offline channels, heterogeneity in the post-adoption mix of spending across channels is likely to translate into systematic differences in profitability across adopter types, even when their total spend is similar.
In summary, our theoretical framework predicts that (i) adopters will increase spending relative to offline-only customers, but (ii) the magnitude and persistence of this effect, as well as the resulting channel mix and profitability, will differ systematically across organic, macro-shock-driven, and promotion-driven adoption pathways, driven by consumer inertia and habit-formation, forward buying around promotions, and the relative profitability of the two channels.
Empirical Setting and Methodology
To test these predictions, we use data from a large pet supplies retailer in Brazil with an extensive network of brick-and-mortar stores and an online presence (website and mobile app). Our data comprise all consumer purchases across all channels from 2019 to 2024. During this period, there were three key events: a macro-environmental shock (COVID-19) and two firm-driven promotions designed to incentivize online shopping. This allows us to observe a variety of adoption pathways.
@@
We construct four groups of adopters of online shopping who previously only shopped at the firm’s physical stores: (1) covid adopters, who started shopping online during the onset of COVID-19; (2) Black Friday adopters, who adopted online shopping due to deep discounts available online during Black Friday; (3) Loyalty Program adopters, who adopted online shopping during the launch of a loyalty program that had to be activated on the retailer’s mobile app; and (4) organic adopters, who started shopping online without any identifiable external stimulus. In addition, we consider a control group of non-adopters or offline-only customers (consumers who did not adopt online shopping during our observation period).
Using these data, we construct consumer-level monthly panel data to quantify revenue, profitability, and the share of spend in the offline and online channels, and use this panel to answer our research questions. Specifically, we measure the Average Treatment Effect on the Treated (ATT) of online adoption on post-adoption behavior for different adoption pathways using a difference-in-differences framework with individual and time fixed effects. Our results should be interpreted as descriptive with appropriate controls; we are not focused on estimating the causal Average Treatment Effect (ATE) of online adoption. Each adoption pathway represents a different type of consumer self-selection, i.e., different types of consumers self-select into adopting online shopping as a result of each stimulus. Our goal is not to rule out selection effects or separate out causal effects from selection effects; rather, we seek to document whether and how different reasons for adopting online shopping shape post-adoption behavior. From a managerial perspective, firms cannot randomize online adoption across consumers; they can only “encourage” adoption through promotional activities. Thus, understanding how different adoption pathways affect post-adoption purchase behavior is of primary interest to managers as opposed to a causal ATE.
Main Findings, Managerial Implications, and Contributions
We now summarize our main findings and their managerial and substantive implications. First, relative to offline-only customers, adopters of online shopping have higher spend and profitability in the post-adoption period, consistent with our predictions and prior evidence on multichannel shoppers. However, the post-adoption behaviors of different adopter types diverge sharply across three key dimensions: total spend, profitability, and the share of offline/online spending. While covid adopters spend similarly to organic adopters post-adoption, promotion-driven adopters (Black Friday adopters and Loyalty Program adopters) spend less. Consistent with our theoretical predictions, this reduced spend is driven largely by forward buying during the promotional period: these customers stock up and then reduce future purchases, which does not occur for covid adopters.
@@
In terms of the share of offline/online spend post-adoption, and consistent with consumer inertia and habit theory, all three adopter groups motivated by external stimuli (covid adopters, Black Friday adopters, and Loyalty Program adopters) have a lower online spend share post-adoption relative to organic adopters. Moreover, covid adopters show slower growth in online spend share over time, consistent with their being older and more set in their habits. Black Friday adopters, by contrast, are younger and show a faster growth rate in online spend share relative to organic adopters, although they still take time to catch up. There is no difference in trends for online spend share between Loyalty Program adopters and organic adopters, likely because these two groups are similar demographically and because the loyalty program involves monthly online activation, reinforcing online shopping habits. Post-adoption profits for the different adopter groups depend on both total spend and the offline/online spend share. We find that despite similar spend, covid adopters are more profitable post-adoption than organic adopters, due to their higher spend share in the offline channel, where margins are higher. In contrast, Black Friday adopters and Loyalty Program adopters are less profitable due to their lower overall spending in the post-adoption period.
Taken together, these findings show that the reason for a consumer’s decision to adopt online shopping significantly shapes long-term multichannel behavior and that the observed patterns are broadly consistent with theories of consumer inertia, age-related habit strength, and forward buying around promotions. Because our mechanisms are grounded in these behavioral forces, the insights are likely to apply more broadly in settings where firms use promotions or external triggers to induce adoption of new channels or digital services.
Our results have important implications for two core managerial tasks: forecasting Customer Lifetime Value (CLV) and evaluating the ROI of promotions. We show that ignoring adoption pathways can lead to substantial bias in incremental CLV forecasts. In our setting, using organic adopters as proxies for promotion-induced adopters overstates the incremental CLV of promotion-induced adopters by roughly 70–80%, and pooling all adopters into a single group similarly misstates the CLV of both promotion-induced and covid adopters. Likewise, breakeven and ROI calculations for promotions can be severely distorted if managers implicitly assume that promotion-induced adopters will behave like organic adopters: in our data, naive breakeven times for online-adoption promotions can be off by factors of three to six. These results imply that firms should (a) explicitly account for channel adoption pathways when estimating incremental CLV from online adoption and (b) set pathway-specific caps on acceptable acquisition and promotional costs.
Overall, our analysis makes both substantive and managerial contributions. Substantively, we extend the multichannel shopping literature by introducing channel adoption pathways as a key dimension of heterogeneity and by documenting how organic, macro-shock-driven, and promotion-driven adopters follow systematically different post-adoption trajectories in spend, channel mix, and profitability. Managerially, we provide a cautionary tale against naively extrapolating the behavior and profitability of one adopter group (e.g., organic adopters) to others and offer a framework for incorporating adoption pathways into CLV forecasting, promotion design, and ROI/breakeven analysis. Firms that explicitly condition on adoption pathways when designing and evaluating interventions to move customers online can better avoid overpaying for low-value adopters and more effectively nurture high-value ones.
@@ 2 Related Literature
Paper Setting Channels Methodology Adoption Pathways
@@
(subscribers only) SMS, Internet, mailing campaign
@@
Our work relates to the empirical literature on how multichannel versus single-channel shopping affects consumers’ purchase behavior. Table 1 summarizes the most closely related papers, and below we highlight how our study differs from and contributes to this literature. For broader reviews, see Neslin et al. (2006); Neslin and Shankar (2009); Verhoef (2012).
From a substantive perspective, our paper contributes in two main ways (see Table 1).
  • $\bullet$
    First, the existing literature overlooks heterogeneity in channel adoption pathways, implicitly assuming that the value of multichannel customers is independent of the motivations underlying their adoption decisions. As Table 1 shows, almost all prior work focuses on organic adopters. Montaguti, Neslin, and Valentini (2016) is an exception, using a mailing campaign to nudge consumers into adopting multiple channels. However, they do not compare organic multichannel customers to those nudged by the campaign, whereas this comparison is central in our paper.
  • $\bullet$
@@
    Second, we study a specific single-channel to multichannel transition that has not been examined before and is particularly relevant for traditional brick-and-mortar retailers such as Walmart, Target, and Costco: moving from shopping only in physical stores to using both physical stores and the retailer’s website/mobile app. Earlier work largely considers transitions from any single channel to some combination of multiple channels (e.g., physical stores, catalogs, phone, mail order, salespeople, mobile apps, websites). These findings may not generalize to the physical-store-to-e-commerce transition because the channels differ in important ways. [^1] Websites and apps provide easy search and comparison between products and between competing retailers, which is not true for catalogs, salespersons, or phone shopping. In many cases, prices online are also lower than in physical stores, which again differs from other channels.